OpenStudy (anonymous):

Solve the equation 1/5 x + 2 = 8 for x. 50 30 40 10

OpenStudy (campbell_st):

subtract 2 from both sides then multiply both sides by 5

OpenStudy (anonymous):

1/5x + 2 = 8 - 2 -2 1/5x = 6 *5 *5 x = 30